The sales representative here tells you they also have two floor plans available, but they only have 38 homes available. Write an equation that illustrates the situation. Use x and y to denote floor plan #1 and floor plan #2 respectively.
For this part I did, x+y=38 but from here I'm not really sure if I know how to do this.
The representative tells you that floor plan #1 sells for $175,000 and floor plan #2 sells for $200,000. She also mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates this situation. Use the same variables you used in part a.
Use elimination to determine how many houses with each floor plan are available. Explain how you arrived at your answer.
I started by making this equation: 175,000x + 200,000y = 7,200,000 I think I am on the right track but I am not sure, word problems and Algebra alone confuse me very much.

Answer by Theo(2978)

you were on the right track.
-----
you let x = number of houses with floor plan 1
you let y = number of houses with floor plan 2
-----
you had x + y = 38
you had 175000*x + 200000*y = 7200000
-----
you need to solve both these equations simultaneously.
-----
you can do it by substitution.
-----
take the first equation and solve for x or y.
either one will do.
I used x = 38-y
-----
substitute for x in the second equation to get:
175000 * (38-y) + 200000 * y = 7200000
this expands to be:
175000 * 38 - 175000 * y + 200000 * y = 7200000
this simplifies to:
6650000 + 25000y = 7200000
subtract 6650000 from both sides of this equation to get:
25000y = 550000
divide both sides by 25000 to get:
y = 22
that makes x = 16 because 22 + 16 = 38
-----